Find the power absorbed by each components, label P1, P2, P3 and P4.


1
Expert's answer
2021-05-05T07:41:12-0400

The easiest way to find the power absorbed by each components is to find it according to one of the following equations:


"P=IV=I^2R=\frac{V^2}{R}."